pooja195 (pooja195):

A teacher is going to give 40 question test worth 100 points.There are two point and four point questions on the test.How many of each type of question is on the test?

Nnesha (nnesha):

same like last one :)

pooja195 (pooja195):

@Nnesha i dunno how ;-;

Nnesha (nnesha):

first what you suppose to find ?? x represent what and y should represent what ?

Nnesha (nnesha):

what are 2 different types ?? :-)

pooja195 (pooja195):

sorry was afk um x=test questions?

OpenStudy (anonymous):

let the questions for 2 points be x and those for 4 points be y. then total questions = x+ y thus, x + y = 40 also the maximum score is 100, so 2x +4y = 100 Solve the two equations for x and y

Nnesha (nnesha):

these are two different types